OpenStudy (anonymous):

The glass for a window is 7.5 feet wide. About how high must a doorway be in order for a contractor to get the glass through the door if the doorway is 3 feet wide?

OpenStudy (dumbcow):

if the diagonal of doorway is 7.5 it will get through --> (7.5)^2 = h^2 + 3^2 solve for h

OpenStudy (anonymous):

so you would use the pythagorean Th?

OpenStudy (anonymous):

≈ 6.87 ft

OpenStudy (anonymous):

How would you go about solving it?

OpenStudy (dumbcow):

get h by itself h^2 = (7.5)^2 - 3^2 take square root h = sqrt(7.5^2 - 3^2)